WHEN IS AN EQUALLY-WEIGHTED DESIGN D-OPTIMAL?

RAMÓN ARDANUY^{1}, J. LÓPEZ-FIDALGO^{1}, PATRICK J. LAYCOCK^{2} AND WENG KEE WONG^{3}

1Departamento de Matemáticas, Pura y Aplicada, Universidad de Salamanca, Spain 2Department of Mathematics, Institute of Science and Technology, University of Manchester, P.O. Box 88, Manchester, M60 1QD, U.K. 3Department of Biostatistics, University of California, Los Angeles, 10833 Le Conte Ave., Los Angeles, CA 90095-1772, U.S.A.

(Received April 7, 1997; revised May 18, 1998)

Abstract. We discuss conditions under which an equally weighted design is D-optimal. The concept of a model being maximally invariant with respect to a set is introduced and is shown to be useful in this work. Several illustrative examples are given, including two biomedical applications using compartmental and segmented models.

Key words and phrases: D-optimum design, equivalence theorem, G-optimum design, information matrix.